Using laws of exponents, simplify and write the answer in exponential form:   (3^4 )^3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of taking power of as power = (a^m)^n= a^mn
(3^4)^3can be written as = (3)^4 × 3
= 3 ^12

Thanks